# Impacts of Data Splitting Strategies on Parameterized Link Prediction Algorithms

### v1

@@ Abstract
Link prediction is a fundamental problem in network science, aiming to infer potential or missing links based on observed network structures. With the increasing adoption of parameterized models, the rigor of evaluation protocols has become critically important. However, a previously common practice of using the test set during hyperparameter tuning has led to human-induced information leakage, thereby inflating the reported model performance. To address this issue, this study introduces a novel evaluation metric, Loss Ratio, which quantitatively measures the extent of performance overestimation. We conduct large-scale experiments on 60 real-world networks across six domains. The results demonstrate that the information leakage leads to an average overestimation about 3.6%, with the bias reaching over 15% for specific algorithms. Meanwhile, heuristic and random-walk-based methods exhibit greater robustness and stability. The analysis uncovers a pervasive information leakage issue in link prediction evaluation and underscores the necessity of adopting standardized data splitting strategies to enable fair and reproducible benchmarking of link prediction models.
@@ 1 Introduction
Link prediction is a fundamental problem in network science, aiming to infer potential or future connections between pairs of nodes based on observed topological information [1, 2, 3, 4, 5, 6, 7]. The practical relevance of link prediction spans a wide range of domains. In social networks, it aids in identifying latent social relationships; in biological networks, it facilitates the discovery of previously unverified protein-protein interactions; and in recommender systems, it enhances the identification of user preferences and improves recommendation quality [8, 9, 10]. Consequently, link prediction has become an essential analytical tool for understanding, modeling, and leveraging the complex relational structures. Over the past few decades, link prediction techniques have undergone substantial evolution. The field has evolved from simple similarity indices based on local structural patterns [11, 8, 12, 13, 14, 15, 16, 17], through parameterized models with enhanced expressive power [18, 19, 20, 21], and more recently toward deep graph representation learning methods [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Research efforts continue to advance link prediction in both theoretical foundations and practical applications. As real-world networks continue to expand in scale, heterogeneity, and temporal dynamics, link prediction algorithms have generally become more complex, with larger model sizes and higher-dimensional parameter spaces [6]. Consequently, the reliability and fairness of performance evaluation procedures have become critically important. Evaluation outcomes not only determine the credibility of empirical comparisons across different methods but also directly influence the methodological trajectory and pace of development in this field. If the evaluation process contains biases or implicit assumptions, algorithmic performance may be inaccurately estimated, leading to misleading conclusions and adverse implications for subsequent research and applications.
@@
In the evaluation of parameterized link prediction algorithms, a critical yet often overlooked issue is the improper use of data splitting strategies. Conventional studies predominantly adopt the two-set split, i.e., the tain–test split, in which the link set is divided into a training set and a test set. The model is first trained on the training subset, and hyperparameter tuning is subsequently conducted using the test set. This strategy has been widely adopted due to the simplicity of implementation [35, 36]. However, the two-set split introduces a fundamental methodological flaw because the model receives feedback from the test set during hyperparameter tuning, which compromises the independence of evaluation. Specifically, in link prediction tasks, researchers typically explore a range of hyperparameter combinations and compute evaluation metrics on the test set, selecting the configuration that yields the best performance (e.g., the highest AUC value). Although this process may seem reasonable, it effectively allows the model to iteratively adjust its hyperparameters based on feedback from the test set [37], thereby inducing a dependency of the learned parameters on the test data. Although the test links are not explicitly used during training, this practice nevertheless violates the basic principle of evaluation independence. The model indirectly acquires structural information from the test set during hyperparameter tuning, resulting in information leakage. In effect, the test set no longer represents truly unseen data; the situation is analogous to having prior access to exam questions before taking a test, granting the model an unwarranted advantage during evaluation. For parameterized algorithms, such leakage leads to inflated test performance, the model appears to perform well because it has been implicitly tuned toward patterns specific to the test set. In contrast, non-parametric algorithms do not involve hyperparameter tuning and thus cannot exploit feedback from the test set, creating an inherent inequity between the two classes of methods [38, 39]. The informational asymmetry compromises the comparability of algorithms and prevents experimental outcomes from accurately reflecting model performance in truly unknown environments, thereby undermining the reliability and reproducibility of link prediction research.
@@
In the field of machine learning, a fundamental principle for preventing information leakage is the strict separation of parameter selection from performance evaluation [39, 40, 41]. Accordingly, when applied to the link prediction task, this principle naturally adopts the three-set splitting strategy, i.e., the train-validation-test split, in which the link set is partitioned into mutually exclusive training set, validation set, and test set. The training set is used for model fitting, the validation set is employed for hyperparameter tuning or early stopping, and the test set is reserved exclusively for the final evaluation after the model is fully specified. By design, the three-set split prevents premature exposure to test data and provides a more faithful assessment of link prediction performance.
@@
However, due to historical inertia and experimental convenience, the two-set split continues to be widely adopted in link prediction research [1, 2, 3, 4, 5, 6, 7]. In many early studies, test set information was mistakenly used during the hyperparameter tuning stage, enabling parameterized algorithms to gain unfair performance advantages and leading to misleading conclusions regarding their true generalization ability on unseen data [1]. Although the three-set split provides an effective solution to this issue, its application in the link prediction domain remains insufficiently widespread. Furthermore, quantitative studies that systematically compare the evaluation discrepancies between the two-set and three-set splits across different algorithmic paradigms and network categories are still lacking.
@@
Existing literature has recognized that data splitting strategies can significantly affect the reliability of model evaluation, yet most discussions are fragmented and limited to specific domains. For example, Roberts et al. [42] pointed out that conventional random splitting in ecological data breaks sample independence, consequently underestimating prediction error, and proposed a structured cross-validation strategy to improve the fairness of generalization assessment. Breit et al. [43] introduced the OpenBioLink benchmarking framework to reveal potential information leakage and unfair comparisons caused by generic data partitioning strategies in link prediction evaluation, and provided a unified and reproducible evaluation protocol. Similarly, Varoquaux and Cheplygina [44] further observed that similar information leakage issues are prevalent in medical machine learning and often lead to inflated evaluation results. Overall, although these studies reveal that improper evaluation procedures may cause biases, they primarily focus on domain-specific scenarios and do not offer a unified quantitative framework applicable across algorithms and network categories.
@@
To analyze the gap between the two-set and the three-set splits from information leakage, we will introduce the metric called Loss Ratio to quantify the performance discrepancy and robustness variation induced by the two-set split, compared to the three-set split, in parameterized link prediction tasks. Thereby, we establish a more equitable and reproducible evaluation standard for link prediction. Many representative algorithms spanning multiple methodological paradigms, including heuristic methods, random-walk-based methods, similarity-transfer methods, and deep learning methods are selected for analysis, including Katz [11], LHN-II [12], Local Path (LP) [15], Local Random Walk (LRW) [16], Superposed Random Walk (SRW) [16], Random Walk with Restart (RWR) [14], TSAA [45], TSCN [45], DeepWalk [24], Graph Convolutional Network (GCN) [28] and Variational Graph Normalized Autoencoder (VGNAE) [33]. This comprehensive comparison across disparate algorithms and networks not only reveals the potential impact of information leakage on the fairness of evaluation but also provides empirical evidence for establishing a more rigorous evaluation framework in the domain of link prediction.
@@ 2 Methods
To characterize the biases introduced by different data splitting strategies, we propose a unified framework that directly quantifies the performance differences caused by test information leakage. This framework compares the performance differences between the two strategies on a common evaluation set, ensuring the fairness. Given a network $G(\mathcal{V},\mathcal{E})$, this framework is designed according to two principle: (1) nested partitioning with a consistent test set, and (2) a unified partition ratio. Specifically, both the two-set and three-set splits are evaluated on an identical test set, ensuring that any observed performance difference can be attributed solely to the partitioning strategy. If we denote $E_{T}$, $E_{V}$, and $E_{P}$ as the training, validation, and test link sets $(\mathcal{E}=E_{T}\cup E_{V}\cup E_{P})$, respectively, and $E_{T}$’ as the training set used in the two-set split. Under the two-set split, the training set $E_{T}^{\prime}$ is first extracted from the link set $\mathcal{E}$ and the remaining links are designated as the unified test set $E_{P}$. Subsequently, we further divide $E_{T}^{\prime}$ into the training set $E_{T}$ and the validation set $E_{V}$, as required by the three-set split, such that $E_{T}^{\prime}=E_{T}\cup E_{V},\quad E_{T}\cap E_{V}=\varnothing,\quad E_{T}\cap E_{P}=\varnothing,\quad E_{V}\cap E_{P}=\varnothing$.
@@
Under the two-set split, the training set $E_{T}^{\prime}$ and the test set $E_{P}$ are jointly used for both model training and hyperparameter tuning. In contrast, under the three-set split, the training set $E_{T}$ is used to learn the model parameters, the validation set $E_{V}$ is employed for hyperparameter selection, and the test set $E_{P}$ is strictly reserved for final evaluation only after the model has been fully determined. To enable a direct comparison between the two strategies, the optimal hyperparameter $\lambda^{\prime}$ obtained by the validation set in the three-set split is projected onto the test set of the two-set split and compared with the performance corresponding to the hyperparameter $\lambda^{*}$ that is obtained directly by the test set (with information leakage). The core procedure is illustrated in Fig. 1.
According to the early studies [46, 49, 48, 47], we adopt the Area Under the ROC Curve (AUC) as the evaluation metric in this work. Then, the Loss Ratio metric is defined as
@@
Heuristic methods mainly rely on the topological information of the network. Katz [11] measures the global connectivity between nodes by summing over all possible paths with an exponential decay factor, thereby capturing long-range structural dependencies. LHN-II [12] can be regarded as a normalized variant of the Katz index, recursively defining node similarity based on the similarity of their neighbors, i.e., if two nodes have similar neighbors, they are also considered similar. LP [15] extends the Common Neighbor by incorporating three-order path information, balancing local accuracy and computational efficiency.
Random-walk-based methods estimate node similarity by simulating information propagation over the network. LRW [16] performs a finite-step random walk starting from a given node and defines similarity according to the probability of visiting another node within a fixed number of steps. SRW [16] aggregates results from multiple walk lengths to emphasize nearby nodes and suppress noise from distant ones. RWR [14] measures global structural proximity based on the expected hitting time in a random walk process, a higher visiting probability between nodes $i$ and $j$ implies greater structural similarity.
Similarity-transfer methods are based on the assumption that similarity between nodes is transitive. TSAA and TSCN [45] extend the Adamic–Adar and Common Neighbor indices, respectively, by introducing a decay factor to aggregate multi-step similarities, thereby capturing indirect structural associations across the network.
Deep learning methods automatically capture dependencies between nodes through graph representation learning. DeepWalk [24] introduces the Skip-Gram model from natural language processing into network representation learning. It performs multiple random walks on the graph to generate node sequences and trains a neural language model to learn low-dimensional embeddings, where structurally similar nodes are mapped close to each other in the embedding space. Graph Neural Networks (GNNs) [28, 30, 29] aggregate neighborhood features through a message-passing mechanism to capture higher-order structural information. In this study, GCN is implemented within the Graph Autoencoder (GAE) framework [28], where graph convolution enables smooth feature propagation and structural integration. VGNAE [33] extends the Variational Graph Autoencoder (VGAE) [28] by introducing normalization constraints and a KL-divergence regularization term. By modeling the uncertainty of latent node representations, VGNAE enhances robustness against noise and alleviates overfitting.
@@ 3 Data Description
This study includes 60 real-world networks [50, 51], covering six categories: Animal Social (13), Biological (11), Informational (8), Social (10), Technological (8), and Transportation (10). The structural statistics of those networks are reported in Table LABEL:DetailofNetwork, where $N=|\mathcal{V}|$ denotes the number of nodes, $M=|\mathcal{E}|$ represents the number of links, $\langle k\rangle$ is the average degree, and $r$ denotes the link density (defined as the ratio of actual links to the theoretical maximum number of links). Overall, those networks vary substantially in both size and density, spanning several orders of magnitude. The diversity ranges from sparse biological and engineering systems to relatively dense social and informational networks, demonstrating that the datasets provide broad coverage and strong representativeness.

### v2

@@ Abstract
Link prediction is a fundamental problem in network science, aiming to infer potential or missing links based on observed network structures. With the increasing adoption of parameterized models, the rigor of evaluation protocols has become critically important. However, a previously common practice of using the test set during hyperparameter tuning has led to human-induced information leakage, thereby inflating the reported model performance. To address this issue, this study introduces a novel evaluation metric, Loss Ratio, which quantitatively measures the extent of performance overestimation. We conduct large-scale experiments on 60 real-world networks across six domains. The results demonstrate that the information leakage leads to an average overestimation of about 3.6%, with the bias reaching over 15% for specific algorithms. Meanwhile, heuristic and random-walk-based methods exhibit greater robustness and stability. The analysis uncovers a pervasive information leakage issue in link prediction evaluation and underscores the necessity of adopting standardized data splitting strategies to enable fair and reproducible benchmarking of link prediction models.
keywords:
Complex Networks , Link Prediction , Data Split Strategies , Evaluation Fairness , Loss Ratio , Information Leakage
journal: Physica A: Statistical Mechanics and its Applications
{highlights}
Revealing information leakage in traditional parameter training for link prediction.
Proposing a method to measure the impact of information leakage.
Demonstrating its significant impact on evaluation results via large-scale experiments.
@@ 1 Introduction
Link prediction is a fundamental problem in network science, aiming to infer potential or future connections between pairs of nodes based on observed topological information lu2011 , wang2015 , martinez2016 , kumar2020 , divakaran2020 , zhou2021 , arrar2024 . The practical relevance of link prediction spans a wide range of domains. In social networks, it aids in identifying latent social relationships; in biological networks, it facilitates the discovery of previously unverified protein-protein interactions; and in recommender systems, it enhances the identification of user preferences and improves recommendation quality adamic2003 , kovacs2019 , lu2012 . Consequently, link prediction has become an essential analytical tool for understanding, modeling, and leveraging the complex relational structures. Over the past few decades, link prediction techniques have undergone substantial evolution. The field has evolved from simple similarity indices based on local structural patterns katz1953 , adamic2003 , leicht2006 , liben2007 , fouss2007 , zhou2009 , liu2010 , cannistraci2013 , through parameterized models with enhanced expressive power clauset2008 , guimera2009 , soundarajan2012 , lu2015 , and more recently toward deep graph representation learning methods kunegis2009 , bordes2013 , perozzi2014 , yang2015 , trouillon2016 , grover2016 , kipf2017 , hamilton2017 , velickovic2018 , zhang2018 , xu2019 , ahn2021 , chamberlain2022 . Research efforts continue to advance link prediction in both theoretical foundations and practical applications. As real-world networks continue to expand in scale, heterogeneity, temporal variability, and diversity of structure formation, link prediction algorithms have generally become more complex, with larger model sizes, higher-dimensional parameter spaces, and more context-dependent zhou2021 , bi2025domain . Consequently, the reliability and fairness of performance evaluation procedures have become critically important. Evaluation outcomes not only determine the credibility of empirical comparisons across different methods but also directly influence the methodological trajectory and pace of development in this field. If the evaluation process contains biases or implicit assumptions, algorithmic performance may be inaccurately estimated, leading to misleading conclusions and adverse implications for subsequent research and applications.
@@
In the evaluation of parameterized link prediction algorithms, a critical yet often overlooked issue is the improper use of data splitting strategies. Conventional studies predominantly adopt the two-set split, i.e., the train–test split, in which the link set is divided into a training set and a test set. The model is first trained on the training subset, and hyperparameter tuning is subsequently conducted using the test set. This strategy has been widely adopted due to the simplicity of implementation kohavi1995 , arlot2010 . However, the two-set split introduces a fundamental methodological flaw because the model receives feedback from the test set during hyperparameter tuning, which compromises the independence of evaluation. Specifically, in link prediction tasks, researchers typically explore a range of hyperparameter combinations and compute evaluation metrics on the test set, selecting the configuration that yields the best performance (e.g., the highest AUC value). Although this process may seem reasonable, it effectively allows the model to iteratively adjust its hyperparameters based on feedback from the test set kaufman2012 , thereby inducing a dependency of the learned parameters on the test data. Although the test links are not explicitly used during training, this practice nevertheless violates the basic principle of evaluation independence. The model indirectly acquires structural information from the test set during hyperparameter tuning, resulting in information leakage. Here, classical data leakage typically refers to sample-level overlap between the training and test data. In contrast, the information leakage discussed in this work specifically denotes an evaluation-induced leakage mechanism, arising from test-dependent hyperparameter selection and manifesting as systematic evaluation bias. In effect, the test set no longer represents truly unseen data; the situation is analogous to having prior access to exam questions before taking a test, granting the model an unwarranted advantage during evaluation. For parameterized algorithms, such leakage leads to inflated test performance, the model appears to perform well because it has been implicitly tuned toward patterns specific to the test set. In contrast, non-parametric algorithms do not involve hyperparameter tuning and thus cannot exploit feedback from the test set, creating an inherent inequity between the two classes of methods dietterich1998 , cawley2010 . The informational asymmetry compromises the comparability of algorithms and prevents experimental outcomes from accurately reflecting model performance in truly unknown environments, thereby undermining the reliability and reproducibility of link prediction research.
@@
In the field of machine learning, a fundamental principle for preventing information leakage is the strict separation of parameter selection from performance evaluation cawley2010 , varma2006 , forman2010 . Accordingly, when applied to the link prediction task, this principle naturally adopts the three-set splitting strategy, i.e., the train-validation-test split, in which the link set is partitioned into mutually exclusive training set, validation set, and test set. The training set is used for model fitting, the validation set is employed for hyperparameter tuning or early stopping, and the test set is reserved exclusively for the final evaluation after the model is fully specified. By design, the three-set split prevents premature exposure to test data and provides a more faithful assessment of link prediction performance.
@@
However, due to historical inertia and experimental convenience, the two-set split continues to be widely adopted in link prediction research lu2011 , wang2015 , martinez2016 , kumar2020 , divakaran2020 , zhou2021 , arrar2024 . In many early studies, test set information was mistakenly used during the hyperparameter tuning stage, enabling parameterized algorithms to gain unfair performance advantages and leading to misleading conclusions regarding their true generalization ability on unseen data lu2011 . Although the three-set split provides an effective solution to this issue, its application in the link prediction domain remains insufficiently widespread. Furthermore, quantitative studies that systematically compare the evaluation discrepancies between the two-set and three-set splits across different algorithmic paradigms and network categories are still lacking.
@@
Existing literature has recognized that data splitting strategies can significantly affect the reliability of model evaluation, yet most discussions are fragmented and limited to specific domains. For example, Roberts et al. roberts2017 pointed out that conventional random splitting in ecological data breaks sample independence, consequently underestimating prediction error, and proposed a structured cross-validation strategy to improve the fairness of generalization assessment. Breit et al. breit2020 introduced the OpenBioLink benchmarking framework to reveal potential information leakage and unfair comparisons caused by generic data partitioning strategies in link prediction evaluation, and provided a unified and reproducible evaluation protocol. Similarly, Varoquaux and Cheplygina varoquaux2022 further observed that similar information leakage issues are prevalent in medical machine learning and often lead to inflated evaluation results. Overall, although these studies reveal that improper evaluation procedures may cause biases, they primarily focus on domain-specific scenarios and do not offer a unified quantitative framework applicable across algorithms and network categories.
@@
To analyze the gap between the two-set and the three-set splits from information leakage, we will introduce the metric called Loss Ratio to quantify the performance discrepancy and robustness variation induced by the two-set split, compared to the three-set split, in parameterized link prediction tasks. Thereby, we establish a more equitable and reproducible evaluation standard for link prediction. Many representative algorithms spanning multiple methodological paradigms, including heuristic methods, random-walk-based methods, similarity-transfer methods, and deep learning methods are selected for analysis, including Katz katz1953 , LHN-II leicht2006 , Local Path (LP) zhou2009 , Local Random Walk (LRW) liu2010 , Superposed Random Walk (SRW) liu2010 , Random Walk with Restart (RWR) fouss2007 , TSAA sun2009 , TSCN sun2009 , DeepWalk perozzi2014 , Graph Convolutional Network (GCN) kipf2017 and Variational Graph Normalized Autoencoder (VGNAE) ahn2021 . This comprehensive comparison across disparate algorithms and networks not only reveals the potential impact of information leakage on the fairness of evaluation but also provides empirical evidence for establishing a more rigorous evaluation framework in the domain of link prediction.
@@ 2 Methods
To characterize the biases introduced by different data splitting strategies, we propose a unified framework that directly quantifies the performance differences caused by test information leakage. This framework compares the performance differences between the two strategies on a common evaluation set, ensuring the fairness. Given a network $G(\mathcal{V},\mathcal{E})$, this framework is designed according to two principle: (1) nested partitioning with a consistent test set, and (2) a unified partition ratio. Specifically, both the two-set and three-set splits are evaluated on an identical test set, ensuring that any observed performance difference can be attributed solely to the partitioning strategy. If we denote $E_{T}$, $E_{V}$, and $E_{P}$ as the training, validation, and test link sets $(\mathcal{E}=E_{T}\cup E_{V}\cup E_{P})$, respectively, and $E_{T}$’ as the training set used in the two-set split. Under the two-set split, the training set $E_{T}^{\prime}$ is first extracted from the link set $\mathcal{E}$ and the remaining links are designated as the unified test set $E_{P}$. Subsequently, we further divide $E_{T}^{\prime}$ into the training set $E_{T}$ and the validation set $E_{V}$, as required by the three-set split, such that $E_{T}^{\prime}=E_{T}\cup E_{V},\quad E_{T}\cap E_{V}=\varnothing,\quad E_{T}\cap E_{P}=\varnothing,\quad E_{V}\cap E_{P}=\varnothing$. In the precedure of link splitting, we preserve the connectivity of the training network $G(\mathcal{V},E_{T})$. More specifically, we will resplit the links if the current splitting results in a disconnected training network.
@@
Under the two-set split, the training set $E_{T}^{\prime}$ and the test set $E_{P}$ are jointly used for both model training and hyperparameter tuning. In contrast, under the three-set split, the training set $E_{T}$ is used to learn the model parameters, the validation set $E_{V}$ is employed for hyperparameter selection, and the test set $E_{P}$ is strictly reserved for final evaluation only after the model has been fully determined. To enable a direct comparison between the two strategies, the optimal hyperparameter $\lambda^{\prime}$ obtained by the validation set in the three-set split is projected onto the test set of the two-set split and compared with the performance corresponding to the hyperparameter $\lambda^{*}$ that is obtained directly by the test set (with information leakage). The core procedure is illustrated in Fig. 1. Importantly, the two-set split is used solely as a contrastive reference for quantifying evaluation bias induced by information leakage, rather than as a recommended or acceptable evaluation procedure.
Given that the primary objective of this work is to analyze the evaluation bias induced by different data splitting strategies, rather than to optimize top-ranked prediction performance in a specific application scenario, we adopt the Area Under the ROC Curve (AUC) as the main evaluation metric, following a series of studies on link prediction metrics zhou2023 , wan2025 , bi2024 , jiao2024 . In the context of link prediction, AUC represents the probability that a randomly selected existing link is assigned a higher prediction score than a randomly selected non-existing link, thereby providing a threshold-free and global assessment of ranking performance. This property makes AUC a stable and neutral basis for comparing evaluation biases across different algorithms and network structures. Then, the Loss Ratio metric is defined as
@@
Heuristic methods mainly rely on the topological information of the network. Katz katz1953 measures the global connectivity between nodes by summing over all possible paths with an exponential decay factor, thereby capturing long-range structural dependencies. LHN-II leicht2006 can be regarded as a normalized variant of the Katz index, recursively defining node similarity based on the similarity of their neighbors, i.e., if two nodes have similar neighbors, they are also considered similar. LP zhou2009 extends the Common Neighbor by incorporating three-order path information, balancing local accuracy and computational efficiency.
Random-walk-based methods estimate node similarity by simulating information propagation over the network. LRW liu2010 performs a finite-step random walk starting from a given node and defines similarity according to the probability of visiting another node within a fixed number of steps. SRW liu2010 aggregates results from multiple walk lengths to emphasize nearby nodes and suppress noise from distant ones. RWR fouss2007 measures global structural proximity based on the expected hitting time in a random walk process, a higher visiting probability between nodes $i$ and $j$ implies greater structural similarity.
Similarity-transfer methods are based on the assumption that similarity between nodes is transitive. TSAA and TSCN sun2009 extend the Adamic–Adar and Common Neighbor indices, respectively, by introducing a decay factor to aggregate multi-step similarities, thereby capturing indirect structural associations across the network.
Deep learning methods automatically capture dependencies between nodes through graph representation learning. DeepWalk perozzi2014 introduces the Skip-Gram model from natural language processing into network representation learning. It performs multiple random walks on the graph to generate node sequences and trains a neural language model to learn low-dimensional embeddings, where structurally similar nodes are mapped close to each other in the embedding space. Graph Neural Networks (GNNs) kipf2017 , velickovic2018 , hamilton2017 aggregate neighborhood features through a message-passing mechanism to capture higher-order structural information. In this study, GCN is implemented within the Graph Autoencoder (GAE) framework kipf2017 , where graph convolution enables smooth feature propagation and structural integration. VGNAE ahn2021 extends the Variational Graph Autoencoder (VGAE) kipf2017 by introducing normalization constraints and a KL-divergence regularization term. By modeling the uncertainty of latent node representations, VGNAE enhances robustness against noise and alleviates overfitting.
@@ 3 Data Description
This study includes 60 real-world networks kunegis2013 , rossi2015 , covering six categories: Animal Social (13), Biological (11), Informational (8), Social (10), Technological (8), and Transportation (10). The structural statistics of those networks are reported in Table LABEL:DetailofNetwork, where $N=|\mathcal{V}|$ denotes the number of nodes, $M=|\mathcal{E}|$ represents the number of links, $\langle k\rangle$ is the average degree, and $r$ denotes the link density (defined as the ratio of actual links to the theoretical maximum number of links). Overall, those networks vary substantially in both size and density, spanning several orders of magnitude. The diversity ranges from sparse biological and engineering systems to relatively dense social and informational networks, demonstrating that the datasets provide broad coverage and strong representativeness.